Faith normally saves £2.50 a week. She decides to save up for a new bicycle and increases her savings such that she now puts a total of £4.50 a week aside. The bicycle costs £300. How much sooner (to the nearest week) will faith be able to purchase her bicycle assuming she begins with no savings?
Answer by greenestamps(13200)



(1) 300/2.50 = 120; 120 weeks if saving 2.50 per week.
300/4.50 = 66 2/3; 66 2/3 weeks if saving 4.50 per week.
difference: 120-66 2/3 = 53 1/3.

She will be able to purchase the bicycle 53 1/3 weeks sooner.


or...

(2) 300/2.50 = 120; 120 weeks if she continued saving 2.50 per week.
4.50/2.50 = 9/5; she decides to save 9/5 as much each week.
It will then take her 5/9 as long to save the required amount.
120*(5/9) = 600/9 = 66 2/3.
120 - 66 2/3 = 53 1/3

Obviously the same answer as by the other method.